OpenStudy (anonymous):

evaluate the expression for x=4 4*(x-2)^2-9 help plz thank you .

Parth (parthkohli):

\( \color{Black}{\Rightarrow 4(4 - 2)^2 - 9 }\) Does that help?

OpenStudy (anonymous):

4*4-9=7

Parth (parthkohli):

Solve for the parentheses first

OpenStudy (lgbasallote):

then the exponent

Parth (parthkohli):

P E M D A S

OpenStudy (anonymous):

it depends on how you look at it... is 2 tge exponent or 2-9?

OpenStudy (lgbasallote):

B O D M A S

Parth (parthkohli):

B O D M A S doesn't have the exponent part but...umm it's ok

OpenStudy (lgbasallote):

i thought we were enumerating acronyms with MAS? darn :/ i was ready for my next word as XMAS

OpenStudy (anonymous):

4*2=8 *2=16=7